A shop keeper bought 7/9 kg of potatoes. He packed them in to packets. Each containing 14/63 kg of potatoes. Haw many packets of potatoes did he make ?

Grade:7

3 Answers

Arun
4 years ago
Dear student
 
number of packets = total weight / weight of one packet =
\frac{7/9}{14/63} = \frac{7* 63}{9*14} = \frac{49}{14} = 3.5\frac{7/9}{14/63} = \frac{7* 63}{9*14} = \frac{49}{14} = 3.5 = 4 packets
 
Hope it helps
 
In case of any difficulty, Please feel free to ask
Satish sahu
4 years ago
Hello Dakshi,
 
A shopkeeper bought potatoes=7/9 kg
Each packet containing potatoes=14/63 kg
Packets of potatoes he made=(7/9)/(14/63)
          =(7×63)/(9×14) 
          =7/2
          =3.5 =4
Since Packet is always full.
Hence packets made are 4.
 
I hope it helps you.
Yashraj
3 years ago
 
 
number of packets = t weight / weight of a sigle  packet
 
no.of packets
=7/9 multiplied by63/14 
so ans is 4